# Adaptive and robust watermark against model extraction attack

### v1

@@ IV-H3 Mixture of unwatermarked data
In a scenario similar to training only a part of watermarked data, attackers might mix all queried data from the victim model with additional unwatermarked source data for training the imitation model. We set $A^{D}=A$ in this case. Our experiments maintain a fixed number of watermarked data while varying the total training data $|D|$, altering $\rho$ to values of 0.1, 0.25, 0.5, and 0.75. We only detect watermarks on queries already known to the victim model’s owner. The results, as illustrated in Fig 10, reveal that increasing the proportion of watermarked data correspondingly elevates the watermark detection in the text generated by the imitation model, without affecting the performance on downstream QA tasks.
@@ V Conclusion
In this paper, we propose PromptShield, a plug-and-play watermarking method for language models designed to combat model extraction attacks. Unlike existing IP protection watermarking methods that modify the original output or distribution of the language model by heuristic rules, PromptShield offers a seamless and effective solution that minimizes distortion to the model’s distribution by utilizing system prompt guidance. PromptShield leverages the self-reminding capabilities inherent in LLM. By integrating user queries with a self-generated watermark system prompt, PromptShield subtly guides the model to incorporate watermark words into its outputs, achieving an end-to-end watermarking process without manipulating the model’s internal probability distribution or altering the model’s output. PromptShield not only maintains the quality of the generated text but also ensures the watermark’s effectiveness and robustness. Extensive experimental results have validated the efficacy of PromptShield. PromptShield has shown remarkable detection sensitivity, even when the watermark detection under interference. Additionally, Its unobtrusiveness and model-agonist nature make it a practical solution for real-world applications, reducing the risk of detection and reverse engineering by potential attackers. We hope PromptShield offers fresh insights for IP protection in large language models. It provides a viable and efficient strategy for model owners to safeguard their IP rights, fostering fairness and innovation in the LMaaS market.
VI Appendix
VI-A Different types of watermarks
Figure 12: The basic workflow of the four types of watermarking. Based on the protection targets, we categorize watermarks into the following four types: model parameter watermarking, content watermarking, dataset watermarking, and IP protection watermarking.
As shown in Figure 12, we classify 4 types of watermarks: conventional model parameter watermarks, content watermarks, dataset watermarks and watermarks against model extraction attack. The last category have been discussed in II-C. Conventional model parameter watermarking methods [36, 37, 38, 39] prove ineffective in safeguarding the intellectual property of models. Usually, these methods insert some backdoor in the parameters of models. But confronted model extraction attack, attackers do not need to understand the internal details of the victim model, and normal query-answer processes are unlikely to trigger specific parameter-based watermarks. Specifically, content watermarks are embedded within a model’s output, enabling identification of whether it originates from a specific model. Most popular model content watermarking [21, 22, 40, 41] fails to effectively protect the intellectual property of models in model extraction attack scenarios. Although some content watermarks can be learned to some extent through model training [71] (except these distribution-preserving ones [22, 41]), the target of content watermarking is to prove whether any text was generated by a special model is not specially designed learnable. Also, after obtaining the content returned by the victim model’s API, attackers’ operations are unpredictable. Attackers can modify the model content and train their own models, making it challenging for content watermarks to persist after modifications and the training process. To the best of our knowledge, dataset watermarking, largely applied to image datasets [72, 73, 74, 75], has yet to specialize in text carriers. The scenario of dataset watermarking shares some similarities with our IP protection situation, but there are significant differences. The fundamental distinction lies in the fact that, in the dataset scenario, the dataset is known and fixed, whereas in our scenario of model extraction attacks, the queries users pose are entirely unknown. Therefore, watermarks in the model extraction attack scenario cannot rely on specific trigger words or phrases, as the nature of user queries is unpredictable. Instead, an adaptive and highly scalable watermark embedding method is needed to handle arbitrary user queries effectively. In addition, using model content watermark can build a dataset watermark as long as the distribution of watermarked texts and no-watermarked generated text’s distribution is distinguishable.
VI-B Workflow of existing watermarking method against model extraction attack
As introduced in the section on related work, we categorize language model IP protection watermarks into two main types: black-box [15] and white-box methods [16, 76]. Black-box methods embed watermarks without needing knowledge of the model’s internal parameters, primarily through post-processing modifications to the original language model’s output text. White-box watermark embedding, on the other hand, requires knowledge of internal generative parameters like logits and modifies the predicted token probability distribution of the original language model according to the specific watermarking rules. The detailed processes of these two methods are illustrated in the Fig 13.
(a) Black-box post-processing watermark. Blue box is the waternarks generation and yellow box is the watermark extraction.
(b) White-box probability distribution adjustment watermark. Blue box is the waternarks generation and yellow box is the watermark extraction.
Figure 13: Workflows of existing language model IP protection watermarking method. The upper subfigure represents black-box post-processing methods and the lower subfigure depicts white-box probability distribution adjustment methods.
VI-C Calculation formulas of BLEU and ROUGE
We calculate the Cumulative BLEU-4 score, which merges precision scores from unigram to 4-gram, as follows:
$BLEU_{4}=BP\cdot\exp\left(\sum_{n=1}^{4}w_{n}\log p_{n}\right),$ (8)
$BP=\begin{cases}1&\text{if }c>r\\ \exp(1-r/c)&\text{if }c\leq r\end{cases}$ (9)
where $p_{n}$ represents the precision for n-grams (n=1 to 4), $w_{n}$ is the weight for each n-gram precision (typically 0.25), $BP$ is the Brevity Penalty for short texts, $c$ is the length of the generated answer, and $r$ is the effective reference corpus length. We also use ROUGE-Lsum $F_{{Lsum}}$ to evaluate task performance, calculated as:
$F_{{Lsum}}=\sum_{i}\frac{(1+\beta^{2})\cdot R_{L_{i}}\cdot P_{L_{i}}}{R_{L_{i}}+\beta^{2}\cdot P_{L_{i}}}.$ (10)
Here, $R_{L_{i}}$ is the recall for the longest common subsequence (LCS) in the $i$-th sentence, $P_{L_{i}}$ is the precision for the LCS in the $i$-th sentence, and $\beta$ is the balance between recall and precision.
VI-D Case study example
Figure 14: Example of the generation of origin LM, no watermark data finetuned LM, Victim LM, and watermark data trained Imitation LM with the same user query.
In the figure 3, we present a case study example that illustrates the responses to the same user query by different models: the base model $M_{O}$, the non-watermarked fine-tuned model $M_{L}$, the victim model $M_{V}$ under extraction attack with our watermark mechanism, and the imitation model $M_{I}$ trained on watermarked data. It is evident that the imitation model displays a learned inclination towards watermark words based on its underlying base model.
VI-E Property of language model self-generated watermark
We did not directly intervene in the choice of watermark words by the language model, but we qualitatively analyzed the preferences of a ChatGPT-scale language model in selecting watermark words for its generated responses. Our analysis began by defining the watermark recall rate $R_{W}$ as the proportion of watermark words in the imitation model’s output to those associated with the given query:
$R_{W}=\frac{\sum_{w\in WM_{i}}\mathbbm{1}_{{y_{i}}}(w)}{|WM_{i}|}.$ (11)
Here, $|WM_{i}|$ denotes the number of watermark words linked to the query $q_{i}$, and $\mathbbm{1}_{y_{i}}(w)$ represents the indicator function that identifies the occurrence of the watermark word $w$ in the generated text $y_{i}$. We observed the characteristics of both high-recall ($R_{W}>0.9$) and low-recall ($R_{W}<0.1$) response corresponding watermark words. We observe that high-recall sentence corresponding watermark words exhibit the following features:
  • Professionalism or uniqueness: Typically specialized terms or uncommon words, like ’biotechnology’ and ’chronos’, which are distinctive and closely related to specific topics.
  • Cultural or linguistic diversity: Words that encompass different cultures or languages, such as ’café au lait’.
  • Specificity and clarity: Terms that are directly related to specific fields or topics.
  • Elongated Structure: Multisyllabic or compound words like ’temporal allocation’.
Low-recall sentence corresponding watermark words, on the other hand, are characterized by:
  • Commonality or frequency of use: More common or general words like ’needs’ and ’creative’, which appear frequently in various texts and are difficult to distinguish.
  • Abstract nature: Words like ’empiricism’ and ’transcendental’ are not commonly used but are too abstract and may appear in multiple contexts, making them hard to identify.
  • Lack of specific domain connection: Words that do not strongly relate to any particular theme or specialized field.
  • Shorter vocabulary: Monosyllabic or simple words that easily blend with other content in the text.
These insights into watermark word characteristics illuminate the nuanced choices a language model makes in self-generating watermarks, shedding light on the complex interplay between watermark words selection and contextual relevance.
VI-F Watermark Intensity analysis
(a) Watermarks count in HC3
(b) Watermarks count in WILD
Figure 15: The influence of watermark words count. The horizontal axis in the graph represents the number of watermark words automatically generated by the protected model. The blue sample points show 2000 output samples of imitation models trained with watermarked data. The red line indicates the average Sentence Watermark Score of the imitation models corresponding to the given watermark word count.
Due to the adaptive nature of our watermarking method, we don’t enforce embedding a fixed number of watermark words in sentences. Instead, the language model intuitively decides the number and generation of watermark words based on the context. As depicted in Fig 15, analysis of the number of watermark words generated $WM$ (all auto-generated without restrictions) indicates a preference of the language model to produce 2-10 watermark words, with two being the most common. We also explored how the number of watermark words correlates with the Sentence Watermark Score output by imitation models. Notably, there’s a general upward trend in the Sentence Watermark Score as the watermark word count increases. However, the range of variation also expands, meaning that the fluctuation range of the scores increases with the word count. This suggests that despite the sparse embedding of watermark words, the language model can still learn watermark characteristics, and more watermark words enhance these characteristics.

### v2

@@ IV-H3 Mixture of unwatermarked data
In a scenario similar to selective training on watermarked data subsets, attackers might mix all queried data from the victim model with additional unwatermarked source data for training the imitation model. We set $A^{D}=A$ in this case. Our experiments maintain a fixed number of watermarked data while varying the total training data $|D|$, altering $\rho$ to values of 0.1, 0.25, 0.5 and 0.75. We only detect watermarks on queries known to the victim model’s owner since the attacker can blend in any unwatermarked data. The results, as illustrated in Fig 10, reveal that increasing the proportion of watermarked data correspondingly elevates the watermark detection in the text generated by the imitation model. Watermarks are detectable in all cases with p-values below $10^{-2}$, without affecting the performance on downstream QA tasks.
@@ V Conclusion
In this paper, we propose ModelShield, a plug-and-play watermarking method for language models designed to combat model extraction attacks. Unlike existing IP protection watermarking methods that modify model’s output or distribution with heuristic rules, ModelShield uses system prompt guidance to minimize distortion seamlessly. Leveraging the self-reminding capabilities of LLMs, ModelShield integrates user queries with a self-generated watermark system prompt, subtly embedding watermark words into outputs without altering the model’s internal probabilities or outputs. This approach maintains text quality while ensuring watermark effectiveness and robustness. Extensive experiments validate ModelShield’s efficacy, showing remarkable detection sensitivity even under interference. Its unobtrusiveness and model-agnostic nature make it practical for real-world applications, reducing detection and reverse engineering risks. Although our method can resist some model extraction attacks, it has limitations. Attackers might use jailbreak prompts to bypass preset system instructions, making it uncertain if our watermarking will be compromised. Defending against jailbreak prompts is complex and extensively studied, so we do not delve into whether system instructions will be executed. We hope ModelShield offers fresh insights for copyright protection of large language models. It provides a viable and efficient strategy for model owners to safeguard their IP rights at a low cost, fostering fairness and innovation in the LMaaS market.